OpenStudy (anonymous):

mammal age of hide that contains 70% of original amt of c-14

OpenStudy (anonymous):

Run that by me again? I feel like a part of the info needed to answer the question, as well as the actual question is missing.

OpenStudy (anonymous):

I think the question is find the age of the mammal if it's hide contains 70% of its Carbon-14, so this is a carbon-dating question, which involves, if I recall, logarithms? Assuming the half-life of Carbon-14 is 5730 years, then the answer is as follows Given \[0.7=1(0.5)^{x/5730}\] Log both sides (with the Common, base 10 log) \[\log(0.7)=(x/5730)\log(0.5)\] Divide both sides by log(0.5) \[0.515=x/5730\] Multiply both sides by 5730 to get... \[x=2948.5\] So your answer /should/ be 2948.5 years. Hope this helps!